# F-95000 Cergy-Pontoise, France

### v1

@@ 6.5 Handling distinct Clauses
In this section, we show how to handle the distinct clause associated with the operator $count$. Recalling that $count({\tt distinct}\,M_{i})$ counts each $M_{i}$-value only once, independently from the number of its occurrences, we consider analytic queries $\mathcal{AQ}$ with no group-by clause, knowing that the presence of a group-by clause does not raise particular difficulties. Given such a query $\mathcal{AQ}:$ select $count({\tt distinct}~{}M_{i})$ from $T$ where $\Gamma$, we let ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$.
Similarly to the general case of the aggregate $sum$, an effective way of computing ${\sf C\_ans}(\mathcal{AQ})$ is currently unknown to the authors. This is so because, according to Algorithm 2, in order to determine $lub$ (respectively $glb$) without counting the same value more than once, every choice of one $M_{i}$-value in each set ${\sf tuples}(\sigma(M_{i}))$ where $\sigma$ is in $\Sigma(\mathcal{AQ})$ (respectively in $\Sigma^{+}(\mathcal{AQ})$) must be analyzed separately. As such a processing clearly requires to compute all possible choices, it is not polynomial, and so not tractable.
Consequently, instead of computing the exact values of $lub$ and $glb$, we propose a tractable approximation of ${\sf C\_ans}(\mathcal{AQ})$. To this end, we first notice that the values to be counted for determining $lub$ are obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma(\mathcal{AQ})$. Therefore, the set resulting from such a choice can not contain $(a)$ more distinct values than there are m-tuples in $\Sigma(\mathcal{AQ})$ and $(b)$ more distinct values than there are distinct values in the union of all sets ${\sf tuples}(\sigma(M_{i}))$ where $\sigma$ is in $\Sigma(\mathcal{AQ})$. This can be formally stated as $(a)$ $lub\leq|\Sigma(\mathcal{AQ})|$ and $(b)$ $lub\leq|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$, which we write as
@@
If $count\_max$ denotes the infimum defined above, we emphasize that Algorithm 2 can be modified so as to compute $count\_max$ in polynomial time. Indeed, it is easy to see that $|\Sigma(\mathcal{AQ})|$ is equal to the value of $max\_ans$ returned by Algorithm 3 in the case of $count$. However, computing the second cardinality requires extra storage and extra computation to ensure that every value in the union is counted only once.
Regarding the approximation of $glb$, we recall from Algorithm 2 that the values to be counted for determining $glb$ are distinct values obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$. Thus, if $\Sigma^{+}(\mathcal{AQ})=\emptyset$, $glb=0$, and otherwise, a trivial lower bound for $glb$ is $1$, because all sets ${\sf tuples}(\sigma(M_{i}))$ for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, are nonempty.
To find a more accurate value of this lower bound, we notice that any choice of one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, concerns at least as many distinct $M_{i}$-values as there are in the union of all ${\sf tuples}(\sigma(M_{i}))$ of cardinality $1$, for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$. Denoting by $\Sigma^{+}_{1}(\mathcal{AQ})$ the set of these m-tuples, that is $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma\in\Sigma^{+}(\mathcal{AQ})~{}|~{}|{\sf tuples}(\sigma(M_{i}))|=1\}$, whenever $\Sigma^{+}(\mathcal{AQ})\neq\emptyset$, we have:
@@
Denoting by $count\_min$ the cardinality shown above, we notice that, as for $count\_max$, Algorithm 3 can be modified so as to compute $count\_min$ in polynomial time.
Another important remark regarding distinct is that if for every $\sigma$ in $m\_Chase(T)$, we have $|{\sf tuples}(\sigma(M_{i}))|=1$ (i.e., the dependency $\mathbf{K}\to M_{i}$ is satisfied), then $\Sigma^{+}_{1}(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$, and:
@@
be an analytic query involving a distinct clause, but no selection condition. In this case, we have $\Sigma(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$ and it can be seen from Figure 2 that the only m-tuples $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ are the first three m-tuples in $m\_Chase(T)$ that we denote by $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$. Thus, $\Sigma(\mathcal{AQ})$ and $\Sigma^{+}(\mathcal{AQ})$ are equal to $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ that are such that ${\sf tuples}(\sigma_{1}(M_{1}))={\sf tuples}(\sigma_{3}(M_{1}))=\{m_{1}\}$ and ${\sf tuples}(\sigma_{2}(M_{1}))=\{m^{\prime}_{1},m^{\prime\prime}_{1}\}$.
Referring to the four repairs of $m\_Chase(T)$ shown in Figure 3, we have ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})=2$ for $i=1,\ldots,4$, implying that ${\sf C\_ans}(\mathcal{AQ})=2$. On the other hand, applying the approximations shown above, we have $|\Sigma(\mathcal{AQ})|=|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=3$ and thus, $count\_max=3$. Moreover, $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma_{1},\sigma_{3}\}$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=1$. Hence, we obtain $count\_min=1$, producing the interval $[1,3]$, which approximates, but is not equal to ${\sf C\_ans}(\mathcal{AQ})$.
Changing $m^{\prime}_{1}$ to $m_{1}$ in $\sigma_{2}$ yields that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})={\sf Ans}(\mathcal{AQ}^{[R_{2}]})=1$ and ${\sf Ans}(\mathcal{AQ}^{[R_{3}]})={\sf Ans}(\mathcal{AQ}^{[R_{4}]})=2$. Thus ${\sf C\_ans}(\mathcal{AQ})=[1,2]$. On the other hand, applying the approximations as above, we have $|\Sigma(\mathcal{AQ})|=3$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=2$ and so, $count\_max=2$. Moreover, $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma_{1},\sigma_{3}\}$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=1$. Hence, we obtain again $count\_min=1$, producing the interval $[1,2]$, which is now equal to ${\sf C\_ans}(\mathcal{AQ})$. $\Box$
@@ 7 Related Work
The problem of consistent query answering in inconsistent databases has motivated considerable research efforts since its introduction in ArenasBC99 . These efforts were first focused on conjunctive and self-join free queries involving no aggregate operators. The case of queries with aggregate operators and possibly with group-by clauses has been introduced in ArenasBCHRS03 and then further investigated in FuxmanFM05 and in KhalfiouiW23 ; KhalfiouiW24 . Focusing on relations whose constraints are unique key-constraints, these approaches rely on earlier work (see Wijsen19 ) characterizing classes of queries that can be rewritten in polynomial time, into SQL queries whose evaluation yields the consistent answer as defined in ArenasBCHRS03 .
In parallel, the issue of consistent conjunctive query answering to queries involving aggregate operators has been addressed in DixitK19 , based on a radically different technique, mainly that of reducing the problem to the well-known SAT problem. Here again, the constraints of interest are key-constraints associated with relations, as in FuxmanFM05 . Comparing our approach with the work in DixitK19 or in FuxmanFM05 , the main differences are the following:
@@
    The functional dependencies in DixitK19 or in FuxmanFM05 do not exactly match those usually considered in star schemas. Indeed, the dependencies $\mathbf{K}\to M_{i}$ are not in the scope of these approaches, contrary to the present approach.
@@
    In our approach, we allow missing values in the input tables, which is not the case in DixitK19 or in FuxmanFM05 .
@@
    Whereas in these approaches conjunctive queries are considered, in our approach, we allow specific disjunctions that we call independent selection conditions. We recall that independent selection conditions generalize conjunctive selection conditions.
  4. 4.
    Dealing with aggregate operators $sum$ and $count({\tt distinct}\,.)$ in our approach may lead to approximate query answers, which is not the case in DixitK19 and in FuxmanFM05 .
Therefore, it turns out that the two approaches are hardly comparable. However, we argue in this section that all results stated in the previous sections also hold when the functional dependencies $\mathbf{K}\to M_{i}$ are not considered. Moreover, in this case, when the aggregate is $count({\tt distinct}\,.)$ on a measure attribute, the consistent answer can be effectively evaluated, instead of being approximated as earlier explained. More precisely, we show that our approach works in the context of DixitK19 or of FuxmanFM05 , assuming that the key of the fact table $F$ consists of all its attributes (namely all key attributes along with all measure attributes).
Indeed, if $FD$ contains no functional dependency of the form $\mathbf{K}\to M_{i}$ then for every $\sigma$ in $m\_Chase(T)$ and every $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is a singleton. Immediate consequences are that, in this context:
  1. 1.
    $FD$ is normalized and acyclic, and so, the results stated in LauS25 hold.
  2. 2.
    Proposition 2(2) is not true: m-tuples in $m\_Chase(T)$ may be distinct and have the same $\mathbf{K}$-value.
  3. 3.
    Step P2 in process (P). is irrelevant and so, should be ignored. Moreover, the second item of Step P3 should be removed because for every $\sigma$ such that $\mathbf{K}M_{i}\subseteq sch(\sigma)$ for $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is reduced to one tuple, and this tuple is consistent. However, it should be emphasized that Proposition 6 still holds in this setting because adapting the proof amounts to discard all arguments related to inconsistencies of tuples involving $\mathbf{K}$-values associated with measure values.
@@
    Despite the first paragraph in the proof of Proposition 9 (mentioning that this proof relies on the fact that $m\_Chase(T)$ cannot contain distinct m-tuples with the same $\mathbf{K}$-value), the proposition still holds. Indeed, as Proposition 6 still applies, for any of the two items in Proposition 9, the modified process (P) generates a repair which satisfies the conditions in the proposition.
@@
    Dealing with aggregate operator $count({\tt distinct}\,.)$ in this context yields the exact consistent answer, instead of an approximation. This is so because in this case all $\sigma(M_{i})$ contain one value, in which case the equalities (1) and (2) in the previous sub-section hold.
An important consequence is that Algorithm 2, and thus Algorithm 3 and Algorithm 4, work in this context, showing that consistent answers to analytic queries can be computed in polynomial time, even when no functional dependency of the form $\mathbf{K}\to M_{i}$ is considered. It can thus be stated that our approach extends approaches from the literature because we allow missing values in the input tables and we can handle some disjunctive selection conditions (namely independent selection conditions involving no key attributes), at the cost of approximate consistent answers when the aggregate operator is $sum$, operating on positive or negative numbers.

### v2

@@ 6.5 Handling distinct Clauses
The clause distinct is used in SQL queries to avoid the processing of duplicates. The most frequent cases in which this key word is used are (1) associated with a select clause, in which case repeated values are not displayed to the user, and (2) associated with aggregation operator countcount, in which case every value is counted only once.
In our approach, it turns out that the first case does not raise specific difficulties, since as usual, this case simply necessitates to check whether a given value has been displayed earlier or not. However, case (2) is much more involved because in order to compute the interval in the consistent answer to a clause of the form select countcount(distinct $M_{i}$)..., it must be known for each repair which $M_{i}$-values have been counted, and this is clearly not tractable. In order to show how to handle the distinct clause associated with the operator countcount, we focus on analytic queries of the form $\mathcal{AQ}:$ select $count({\tt distinct}~M_{i})$ from $T$ where $\Gamma$, because the presence of a group-by clause does not raise any particular difficulties.
Similarly to the general case of the aggregate sumsum, an effective way of computing the interval $[glb,lub]$ of ${\sf C\_ans}(\mathcal{AQ})$ is currently unknown to the authors. This is so because, as mentioned above, repairs must be considered separately in order to compute lublub and glbglb, which cannot be done when running Algorithm 3. Consequently, instead of computing the exact values of lublub and glbglb, we propose a tractable approximation of ${\sf C\_ans}(\mathcal{AQ})$, easily implemented based on Algorithm 3 and Algorithm 4.
To this end, we first notice that the values to be counted for determining lublub are obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma(\mathcal{AQ})$. Therefore, the set resulting from such a choice can not contain $(a)$ more distinct values than the number of m-tuples in $\Sigma(\mathcal{AQ})$ and $(b)$ more distinct values than the number of distinct values in the union of all sets ${\sf tuples}(\sigma(M_{i}))$ where $\sigma$ is in $\Sigma(\mathcal{AQ})$. This can be formally stated as $(a)$ $lub\leq|\Sigma(\mathcal{AQ})|$ and $(b)$ $lub\leq|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$, which we write as
@@
If count_maxcount\_max denotes the infimum defined above, we emphasize that Algorithm 3 can be modified so as to compute count_maxcount\_max. Indeed, it is easy to see that $|\Sigma(\mathcal{AQ})|$ is equal to the value of proc_maxproc\_max returned by Algorithm 3 in the case of countcount. Moreover, the computation of $|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$ can be implemented in Algorithm 3 in a similar way, at the cost of an extra storage to ensure that every value in the union is counted only once. Then, once the two cardinalities are known, count_maxcount\_max can easily be computed and returned by modifying Algorithm 4 accordingly.
Regarding the approximation of glbglb, we recall from Algorithm 3 that the values to be counted for determining glbglb are distinct values obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$. Thus, if $\Sigma^{+}(\mathcal{AQ})=\emptyset$, $glb=0$, and otherwise, a trivial lower bound for glbglb is $1$, because all sets ${\sf tuples}(\sigma(M_{i}))$ for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, are nonempty.
To find a more accurate lower bound, we notice that any choice of one value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, concerns at least as many distinct $M_{i}$-values as there are in the union of all ${\sf tuples}(\sigma(M_{i}))$ of cardinality 11, for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$. Denoting by $\Sigma^{+}_{1}(\mathcal{AQ})$ the set of these m-tuples, that is $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma\in\Sigma^{+}(\mathcal{AQ})~|~|{\sf tuples}(\sigma(M_{i}))|=1\}$, whenever $\Sigma^{+}(\mathcal{AQ})\neq\emptyset$, we have:
@@
Denoting by count_mincount\_min the cardinality shown above, we notice that, as for count_maxcount\_max, Algorithm 3 and Algorithm 4 can be modified so as to compute count_mincount\_min.
Another important remark regarding distinct is that if for every $\sigma$ in m_Chase(T)m\_Chase(T), we have $|{\sf tuples}(\sigma(M_{i}))|=1$ (i.e., the dependency $\mathbb{K}\to M_{i}$ is satisfied), then $\Sigma^{+}_{1}(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$, and:
@@
be an analytic query involving a distinct clause, but no selection condition. In this case, we have $\Sigma(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$ and it can be seen from Figure 2 that the only m-tuples $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ are the first three m-tuples in m_Chase(T)m\_Chase(T) that we denote by $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$. Thus we have $\Sigma(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})=\{\sigma_{1},\sigma_{2},\sigma_{3}\}$, along with ${\sf tuples}(\sigma_{1}(M_{1}))={\sf tuples}(\sigma_{3}(M_{1}))=\{m_{1}\}$ and ${\sf tuples}(\sigma_{2}(M_{1}))=\{m^{\prime}_{1},m^{\prime\prime}_{1}\}$.
Referring to the four repairs of m_Chase(T)m\_Chase(T) shown in Figure 4, we have ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})=2$ for $i=1,\ldots,4$, implying that ${\sf C\_ans}(\mathcal{AQ})=2$. On the other hand, applying the approximations shown above, we have $|\Sigma(\mathcal{AQ})|=|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=3$ and thus, count_max=3count\_max=3. Moreover, $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma_{1},\sigma_{3}\}$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=1$. Hence, we obtain count_min=1count\_min=1, producing the interval $[1,3]$, which approximates, but is not equal to ${\sf C\_ans}(\mathcal{AQ})$.
Changing $m^{\prime}_{1}$ to $m_{1}$ in $\sigma_{2}$ yields that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})={\sf Ans}(\mathcal{AQ}^{[R_{2}]})=1$ and ${\sf Ans}(\mathcal{AQ}^{[R_{3}]})={\sf Ans}(\mathcal{AQ}^{[R_{4}]})=2$. Thus ${\sf C\_ans}(\mathcal{AQ})=[1,2]$. On the other hand, applying the approximations as above, we have $|\Sigma(\mathcal{AQ})|=3$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=2$ and so, count_max=2count\_max=2. Moreover, $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma_{1},\sigma_{3}\}$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=1$. Hence, we obtain again count_min=1count\_min=1, producing the interval $[1,2]$, which is now equal to ${\sf C\_ans}(\mathcal{AQ})$. $\Box$
@@ 7 Related Work
The problem of consistent query answering in inconsistent databases has motivated considerable research efforts since its introduction in ArenasBC99 . These efforts were first focused on conjunctive and self-join free queries involving no aggregate operators. The case of queries with aggregate operators and possibly with group-by clauses has been introduced in ArenasBCHRS03 and then further investigated in FuxmanFM05 . Focusing on relations whose constraints are unique key-constraints, these approaches rely on earlier work (see Wijsen19 ) characterizing classes of queries that can be rewritten in polynomial time, into SQL queries whose evaluation yields the consistent answer as defined in ArenasBCHRS03 . In parallel, the issue of consistent conjunctive query answering to queries involving or not aggregate operators has been addressed in DixitK19 ; DixitK22 , based on a radically different technique, mainly that of reducing the problem to the well-known SAT problem.
By relying on $Chase$ semantics, our approach does not strictly follow the same semantics as those cited above. Indeed, as argued in JIIS ; LauS25 and illustrated in Example 2, in our approach, database consistency is not checked against the database itself but rather, against its ‘chased’ version. It follows that, as explained in Section 4.1, the notion of repair in our approach does not coincide with the usual one in the literature, where a repair is a maximal and consistent subset of the tuples in the database. It follows that, although our definitions of consistent answer to a query, whether standard or analytic, follow the usual definitions (i.e., equal to the intersection of the answers to the query in every repair for standard queries, and equal to the interval containing all aggregate values found in the repairs for analytic queries), our semantics to consistent query answering does not coincide with those in traditional approaches. Nevertheless, since in the present work we concentrate on databases operating over a star schema, our approach is a particular case of databases whose relations have a unique primary key as in DixitK19 ; DixitK22 ; FuxmanFM05 ; KhalfiouiW23 ; KhalfiouiW24 . Moreover, it has been argued (see Proposition 4 in Section 4.2) that, in this case, our notion of repair can be seen as the usual notion of repair applied to the set of all true tuples (which is super-set of the tuples actually stored in the database). We also notice that in DixitK19 ; DixitK22 , and contrary to our approach, constraints other than key dependencies are considered since these approaches also consider denial constraints, which are known to generalize functional dependencies.
[…]
Another important specificity of queries in our approach, whether standard queries or analytic, is that their expression involves a single table (in fact the m-table m_Chase(T)m\_Chase(T)), which can be seen as an extension of the lossless join of all tables in the data warehouse. As a consequence, using the terminology in FuxmanFM05 , standard queries are tree-queries, whose join graph is the tree having the fact table $F$ as its root and all dimension tables $D_{i}$ as its leaves. Hence, as stated by Theorem 1 in FuxmanFM05 , all standard queries in our approach, are SQL-rewritable, meaning that each query $Q$ can be rewritten into an SQL query $Q^{\prime}$ whose answer is the consistent answer to $Q$. The case of analytic queries is however less clear. Indeed, although a similar result has been stated in FuxmanFM05 , it is argued in KhalfiouiW24 that the SQL rewriting for analytic queries involving the aggregate sumsum is flawed, because this aggregate operator is not monotonic. In fact, a consequence of the separation Theorem 1.1 in KhalfiouiW24 is that for any analytic query in our approach whose aggregate operator is monotonic and associative, it is possible to generate two SQL queries, each of them computing a bound of the consistent answer.
The proof of this flaw, that can be found in DixitK22 or in KhalfiouiW24 , relies on a database schema expressed in our approach as follows: the database has three relations $D_{1}$, $D_{2}$ and $F$, respectively, defined over attributes sets $K_{1}A_{1}$, $K_{2}A_{2}$ and $K_{1}K_{2}M_{1}$, over which the two functional dependencies $K_{1}\to A_{1}$ and $K_{2}\to A_{2}$ are assumed. That is, $K_{1}$, respectively $K_{2}$, is the key of $D_{1}$, respectively $D_{2}$, and the key of $F$ is $K_{1}K_{2}M_{1}$ (i.e., $F$ has no non-key attributes). Hence, this schema can be seen as a star schema, whose fact table has no embedded functional dependency. In this context, it is shown in DixitK22 and in KhalfiouiW24 that computing the consistent answer to the analytic query $\mathcal{AQ}:$ select $sum(M_{1})$ from $T$ where $A_{1}=a_{1}$ and $A_{2}=a_{2}$ (where $a_{1}$ and $a_{2}$ are constants respectively in $dom(A_{1})$ and $dom(A_{2})$) is NP-hard.
Although the above case does not follow the context of our work, since the functional dependencies $\mathbb{K}\to M_{i}$ are missing, we argue that all results stated earlier also hold in this case. Indeed, if $FD$ contains no functional dependency of the form $\mathbb{K}\to M_{i}$ then for every $\sigma$ in m_Chase(T)m\_Chase(T) and every $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is a singleton. Immediate consequences are that, in this context:
@@
    FDFD is normalized and acyclic, and so, the results stated in LauS25 hold.
@@
    Proposition 2(2) is not true: m-tuples in m_Chase(T)m\_Chase(T) may be distinct and have the same $\mathbb{K}$-value.
@@
    Step P2 in process (P) is irrelevant and so, should be ignored. Moreover, the second item of Step P3 should be removed because for every $\sigma$ such that $\mathbb{K}M_{i}\subseteq sch(\sigma)$ for $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is reduced to one tuple, and this tuple is consistent. However, it should be emphasized that Proposition 7 still holds in this setting because adapting the proof amounts to discard all arguments related to inconsistencies of tuples involving $\mathbb{K}$-values associated with measure values.
@@
    Despite the first paragraph in the proof of Proposition 10 (mentioning that this proof relies on the fact that m_Chase(T)m\_Chase(T) cannot contain distinct m-tuples with the same $\mathbb{K}$-value), the proposition still holds. Indeed, as Proposition 7 still applies, for any of the two items in Proposition 10, the modified process (P) generates a repair which satisfies the conditions in the proposition.
@@
    Dealing with aggregate operator $count({\tt distinct}\,.)$ in this context yields the exact consistent answer, instead of an approximation. This is so because in this case all $\sigma(M_{i})$ contain one value, implying that the equalities (1) and (2) in the previous sub-section hold.
An important consequence is that Algorithm 3, and thus Algorithm 4 and Algorithm 5, work in this context, showing that consistent answers to analytic queries can be computed in quasi-linear time, even when no functional dependency of the form $\mathbb{K}\to M_{i}$ is considered. As a consequence, based on the results in DixitK22 or in KhalfiouiW24 , determining whether analytic queries in a star schema according to our approach is SQL-rewritable, is an open issue to be investigated.
On the other hand, this also shows that our approach extends approaches from the literature, in the sense that we allow missing values in the input tables and we can handle some disjunctive selection conditions (namely independent selection conditions involving no key attributes), at the cost of approximate consistent answers when the aggregate operator is sumsum, operating on positive or negative numbers, or when the aggregate operator is of the form countcount(distinct .).
To sum up the above remarks, we recall that our approach works according to different semantics than in existing works dealing with consistent answering. Although this prevents from accurate comparisons, we emphasize the following two main important features of our work:
  1. 1.
    Contrary to any other approach, ours supports the presence of missing values in the data warehouse, at the cost of approximate consistent answers when the aggregate operator is sumsum, operating on positive or negative numbers or when the aggregate operator is of the form countcount(distinct .).
  2. 2.
    Our complexity results outperform those from the literature, since we have shown that consistent answers are computed in quasi-linear time, i.e., in $\mathcal{O}(W\cdot\log(W))$, where $W$ is the size of the data warehouse.